OpenStudy (raffle_snaffle):

Davis is saving money for a bicycle trip around the world. He would like to have $23415 saved in 5 years. To reach his goal, how much does Davis need to deposit today in an account paying 7% per year compounded annually?

OpenStudy (raffle_snaffle):

F = p(1+i*n) = 23415*(1+(0.07)(5)) = 31610.25 31610.10-23415 = 8195.25 ?

OpenStudy (raffle_snaffle):

I want to know how to set this up

OpenStudy (r.hamza17):

As in the question maturity amount is given which is A=$23415, the time period n=5 is also given and the rate of interest r=7% is given. Now y using the formula for finding the present amount, A=P(1+r)^n Now put the values to get the answer

OpenStudy (r.hamza17):

This is how you set this up

OpenStudy (raffle_snaffle):

This is compound interest?

OpenStudy (r.hamza17):

Yup this is for the compound interest

OpenStudy (raffle_snaffle):

Okay thank you.

OpenStudy (r.hamza17):

Amigo